OpenStudy (anonymous):

x^2+11x=-22 solve by factoring medal will be given!

OpenStudy (anonymous):

-11+(sqrt33)/2, -11-(sqrt33)/2

OpenStudy (moongazer):

2(2x+11x)=-22 2\2(2x+11x)=-22/2 2x+11x=-11 12x/12=-11/12 x=-11/12

OpenStudy (anonymous):

Wait, wha...? Where'd the x^2 go? You're missing like........... Alot

OpenStudy (moongazer):

oooops.... wrong equation^_^

OpenStudy (anonymous):

haha :)

OpenStudy (anonymous):

(x-11)(x-2)

OpenStudy (moongazer):

sorry for that

OpenStudy (anonymous):

No, it's not (x-11)(x-2) FOIL it and it doesn't come out the same

OpenStudy (anonymous):

so the factor is by grouding (x^2+11x)+25 then you x(x+11)+25 that is your answer

OpenStudy (anonymous):

x^2+11x+22=0 x= [-11+(121-88)^1/2]/2=[-11+(33)^1/2]/2 and one more root is [-11-(33)^1/2]/2